Can someone please show me the step by step process in solving this type of problem? Thanks a bunch.
Terri Hoelker invested some money at 3.5% simple interest, and $5000 more than 3 times this amount at 4%. She earned $1440 in interest. How much did she invest at each rate?

Answer by ptaylor(2198)


I AM ASSUMING THAT SHE INVESTED HER MONEY FOR 1 YEAR
Let x= amount invested at 3.5%
Then 3x+$5000=amount invested at 4%
INTEREST(I)=PRINCIPAL (P) TIMES RATE(R) TIMES TIME(T) or I=PRT
So our equation to solve is:
0.035x+0.04(3x+5000)=1440 get rid of parens
0.035x+0.12x+200=1440 subtract 200 from both sides
0.155x=1440-200
0.155x=1240 divide both sides by 0.155
x=$8000--------------------------amount invested at 3.5%
3x+5000=24000+5000=29000----------------------amount invested at 4%
CK
0.035(8000)+0.04(29000)=1440
280+1160=1440
1440=1440

Hope this helps----ptaylor